OpenStudy (anonymous):

verify the identity: cos(2x) / sin(x) cos(x) = cot(x) - tan(x)

OpenStudy (anonymous):

Working with the right sides: \[\cot x-\tan x=\frac{\cos x}{\sin x}-\frac{\sin x}{\cos x}\] Find a common denominator and simplify. Also, use the fact that \(\cos(2x)=\cos^2x-\sin^2x\).